# Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL

### v3

@@ 5.3. Pipelined Shared Accesses
As a result, we designed a mode where we temporarily change the speculation model to allow shared memory burst accesses (as pipelining is a form of speculation). Nevertheless, we don’t want to do so while enabling any type of speculation. Indeed, we don’t want our program to speculatively jump to a spectre-like gadget and transmit secrets to our attacker. To avoid this, we tweak the speculation model to allow pipelined accesses to shared memory but disable the use and training of any predictors that could result in the speculative execution of rogue code. Specifically, we disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS). Effectively, that guarantees that the hardware will always predict that the next program counter to be “$pc+4$” and avoid speculative execution of arbitrary code. We call this straight-line speculation.
Modifying the speculative model also modifies the hardware software contract. The programmer now needs to ensure that when this mode is enabled, no speculative transmitter can be reached by the program when doing straight-line speculation. Static analysis can be done manually or by reusing many existing tools: e.g. oo7(wang2019oo7, ), Kasper(johannesmeyer2022kasper, ), Kleespectre(wang2020kleespectre, ), Spectector(guarnieri2020spectector, ). Note that the usual critique and limitation of these tools are that 1) the code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model. Interestingly, both of those limitations are not concerns in the context of Citadel’s burst mode because the code snippets we analyze are small (e.g. boiling down to a memcpy function for instance) and the speculation model is well-defined, simple and enforced (only speculates PC+4). Hence, these works are well-suited to our needs. We consider adapting these tools to our more simple setup as future work. We add two new configurations in our software-exposed configuration registers to enable and disable this “burst” mode. The right way to use this mode is, like in our case, during burst accesses to shared memory (e.g. a memcpy function moving data from public memory to private enclave memory). The programmer only needs to identify these series of burst access and guard them with instructions activating and deactivating burst mode. When doing so on our cryptographic benchmark, we observe that overhead falls from $19.4\%$ to $2.7\%$. More results are presented in the evaluation section.
5.4. Relaxed Flushing Policies
Because we build the entire software infrastructure required to run enclaves, we were able to observe that some security domains, especially the enclave and some parts of the mini-SM, do not have secrets to keep from each other. Specifically, the mini-SM never manipulates sensitive data. As a result, the mini-SM does not need to flush microarchitectural state when context switching to enclaves. This reduces the cost of mini-SM call and opens the door to the use of frequent software or timer interrupts. To disable flushing selectively, we modify the hardware to deactivate systematic flushing and exposed to the SM a register to trigger the flush. We write the SM to trigger the flushing mechanism on required context switches, but not on most traps and SM-calls from the enclave. We evaluate the performance implication of our mechanism in Section 6.6.
@@ 6. Evaluation
6.1. Experimental Setup
Frontend
2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
Reorder Buffer
64 entries, 2-way insert/commit
Reservation Stations
2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
L1 (I/D)
32KB, 8 ways, 8 outstanding requests
Load-Store Unit
24 LdQ entries, 14 StQ entries, 4 SB entries
LLC (L2)
1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
Memory
120 cycles latency (24 outstanding requests)
@@ 24 LdQ entries, 14 StQ entries, 4 SB entries
Citadel’s processor is a modified out-of-order 2-core based on Riscy-OOO (zhang2018composable, ) and MI6 (bourgeat2019mi6, ). The parameters used for the cores and the memory subsystem are given in Figure 5. Citadel runs on an AWS EC2 F1 instance, equipped with a Virtex UltraScale+ FPGA. On this FPGA board, Citadel can be clocked at $30$ MHz, which is similar to the baseline design. We use the RISC-V tool chain with GCC 12.2.0. We compile all software with the -O3 flag. We run Linux Kernel 6.2.0. Unless mentioned differently, all performance numbers are collected by running the benchmark on the FPGA using performance counters and an average of ten runs.
6.2. Implementation Data and TCB Breakdown
Component
LOC Size
Processor (bourgeat2019mi6, ) (+ modification)
$\sim 60$ KLOC (+ 250 ) N/A
Bootloader
5236 55KB
Security Monitor
9287 97KB
Mini-SM
2301 5.6KB
Linux
$\sim 20$ MLOC 157MB
SM Kernel Mod.
507 505KB
TCB Software
14523 152KB
@@ 6.2. Implementation Data and TCB Breakdown
6.3. Evaluating the End-to-End Platform
MicroPython Runtime We do a minimal port of MicroPython (microphython, ) inside one of our enclaves. We start with the bare-metal ARM implementation. We modify the configuration files and the linker script so our binary can run in the virtual memory layout provided by our enclave. We only need to implement a few functions by hand. For instance, we redirect the console to print to shared memory and need to write a bit of assembly for the setjmp and longjmp functions used for exception handling. We also add a cryptographic library, so the enclave can sign the console output using its own private keys recovered through the remote attestation mechanism. A hash of the standard output is extended on each character print and signed on the enclave exit. We cross-compile MicroPython using the RISC-V tool-chain, build without standard library, and generate position independent code. In total, we need to manually modify 128 LOC for the port and add the 5K LOC cryptographic library. The final compiled binary measures 180KB. With the enclave’s code added, the TCB is now 65,174 LOC, most of it (70%) being MicroPython.
Figure 6. MicroPython Enclave Execution Breakdown.
We use a simple primality test as a benchmark where a number is found to be prime if none of the numbers smaller than it can divide it. We test the primality of an arbitrary small prime number: $19,531$. The tests take an average of $4.1\times 10^{7}$ cycles to execute, with 51% of the time spent within the Python program. This means the enclave and attestation mechanism add 94% overhead for this small computation. We show a visual breakdown of the running time in Figure 6. We note that most of this overhead comes from hashing enclave code (25% of the total running time) but also generating the enclave keys and signing the certificate (10% of the running time). These two numbers are especially high since this code is run non-speculatively inside the SM. On another hand, hashing and signing the output inside of enclave takes a negligible amount of time (less than 1% of the total running time). Because this cost is fixed, it could definitely be amortized for bigger applications.
Securing a Cryptographic Library We write a wrapper around an ED25519 RISC-V library (ed25519github, ). The wrapper makes it possible to keep the keys used by the library inside of the enclave and for the library functions to be accessed in a Remote Procedure Call style (RPC). We add a queue to shared memory, so untrusted applications can send requests to the library to generate keys (that will stay in enclave memory), or to sign messages using previously generated keys. In total, we need to implement 334 LOC of C, which is to be compared to the 4,514 LOC of the crypto library (7% overhead). With the enclave’s code added, the total TCB is now 19,231 LOC. The final crypto-enclave binary is 71KB.
We compare performance between three scenarios. In each of these scenarios, we measure the number of cycles required to sign randomly generated messages. We split 1MB of randomly generated data into $\approx 3000$ packets of various lengths following the arbitrary Internet Mix distribution (imix, ). In the first setup, which corresponds to our insecure baseline, we do not use any enclave as we directly link the cryptographic library to our user-level application. In the second scenario, called “multi-core”, the enclave constantly runs on a dedicated core and programs running on other cores can forward requests through the shared memory mechanism. In the third scenario, termed “re-entry”, the enclave is re-entered every single time an application calls the cryptographic library on the same core as that of the application. We see very little difference (2.9%) between the first two setups, and a more important 27.6% slowdown between the first and third scenario. Indeed, our cryptographic algorithm is compute-heavy, making the communication overhead negligible in the second scenario. Nevertheless, context-switching in and out of an enclave is expensive as it requires flushing the microarchitectural state, which makes the slowdown more important in the re-entry scenario. These two secure solutions highlight an interesting tradeoff between resource occupation (one core and a DRAM region vs. only a DRAM region) and performance.
Private Inference We evaluate the following scenario where a remote user wants to do perform some private inference task on a neural network evaluated on an untrusted server. We pace a simple neural network model inside of one of our enclaves. We use onnx2c (onnx2cgithub, ) to convert a ONNX file describing a handwritten digit recognition model (trained on MNIST) to standard C code and run private inference inside of our enclave. The user uses our remote attestation mechanism to verify the enclave identity and establish a shared secret with it. The user is then able to send 1024 encrypted requests to the enclave that are passed along to the enclave by the untrusted OS. Upon reception, the enclave decrypts the ciphertext and performs the inference task on the private input. It then sends back an encrypted output to the remote user. In total, we need to implement 227 lines of straightforward C, which is to be compared to the 1536 LOC of the neural network and the 410 LOC of the AES library and the 4386 LOC of ED25519 we now need to include inside of the enclave for encryption and key exchange. With the enclave code added, the TCB is now 21K LOC and the final enclave binary measures 54KB.
We now compare the performance between our setup and an insecure baseline where the model is directly evaluated outside of an enclave and where the LLC is not partitioned. We do consider that incoming traffic from the remote user is encrypted and needs to be decrypted first. In both setups, runtime is split 55% and 45% between AES operations and inference. We see negligible (0.76%) overhead for running a model inside of the enclave, with all of it coming from the necessity to copy the ciphertext in enclave private memory before performing decryption. This is not surprising as the rest of the computation happens in enclave private memory and should not suffer from any slowdown. This doesn’t account for the time required to set up the enclave, which is a one-time cost but represents 23.3% of the cost of performing 1024 inference.
@@ 6.4. Evaluating Secure Shared Memory
We first demonstrate that naive public memory sharing is vulnerable to transient execution attacks by mounting an attack on an insecure version of Citadel where speculation on shared memory is allowed. The attack uses the LLC as a side channel and use a specter gadget located in enclave code to leak arbitrary enclave memory to untrusted code ruining on another core. Anecdotally, we were not able to successfully mount the attack on our secure prototype.
Table 2. Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave. Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program. We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).
@@
We then evaluate the perfromance overhead of our comunication mechanism. We use a similar setup as Section 6 but only measure time spent computing the cryptographic function. The enclave performs $3000$ cryptographic operations (signature or cryptographic hash) over random inputs provided by the OS. We compare three setups. First, our final design with burst mode enabled around shared memory burst sequences (Citadel). Then Citadel without annotating code for shared memory burst sequences (- Burst). Finally, another (insecure) configuration where we annotate all the code to be run in burst mode (++ Burst). Our results are presented in Table 6.4. We report the runtime normalized to the insecure baseline (directly linked library and no LLC partitioning). We observe an overhead of 2.7% on the signature benchmark and 46.6% for the hashing benchmark. This is because the hashing application accesses shared memory much more frequently compared to the signing application that is more computationally bounded. The hashing function is more affected by the loss of the L1 cache. This difference between the benchmarks is also highlighted by our other results, where not annotating the code for burst shared accesses has an $397.3\%$ overhead for the hash application and a $19.4\%$ for the signing applicaiton. These results also demonstrate the efficiency of our burst mechanism when used properly.
We also investigate the performance of using this mode on the whole enclave code instead of the limited code segment when the enclave communicates. On our benchmarks, we observe that the performance overhead goes up to $65.4\%$ for the signing benchmark and $395.9\%$ for the hashing benchmark. We investigate the origin of this overhead and realize that miss speculation due to the deactivation of the RAS generates a lot of random load that pollutes the TLB, triggers many page walks and terribly slows down the virtual memory subsystem. This is due to many small functions being linked one after the other in the binary. Some of these functions start with loads using generic registers as addresses (for instance when accessing a pointer passed as an argument to the function). When encountering a ret instruction, the processor will first speculate that the following instructions should be executed. In this speculative window, it will start running the instructions of the function linked right after and perform loads or stores to random addresses from registers containing garbage. We confirm this hypothesis by instrumenting the binary and observing that when the RAS is reactivated (but not other predictors), the number of TLB misses, page walks and the overall execution time go down significantly. As a result, we recommend a hybrid approach where the burst mode is only used when the enclave wants to communicate.

### v4

@@ 4.5. Burst Mode: Pipelined Shared Memory Accesses
As a result, we designed a complementary mode where we temporarily change the execution mode to allow shared memory burst accesses, but without enabling any type of attacker-controlled speculation. For instance, we don’t want our program to speculatively jump to a Spectre-like gadget and transmit secrets to our attacker. To avoid this, we restore speculative accesses to shared memory, but disable the use and training of all predictors, besides “pc+4”. Specifically, we modify the Fetch pipeline in order to disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS) (see Figure 6). This is commonly referred to as straight-line speculation (straightlinespecARM, ; guarnieri2021hardware, ). Note that, for security, we still need to bypass the L1-D (see Section 6).
Proper Usage The switch between SafeSpec and Burst mode is triggered by writing to a specific Control Status Register (CSRs). Note that CSR reads and writes cannot be performed speculatively, and the corresponding instructions effectively act as speculation barriers by flushing the processor pipeline. The right way to use this mode is on small code snippets that perform back-to-back accesses to shared memory, as shown in Figure 7 or Listing 2 (full code and assembly in supplementary material). The programmer or compiler only needs to identify these series of in-lined burst accesses and guard them with the CSR write instructions activating and deactivating burst mode, effectively isolating the code snippets. Modifying the execution mode also modifies the requirements on the software. In Figure 7, we highlight how the two mechanisms change the (speculative) control flow for the same assembly snippet from memcpy. The programmer now needs to ensure that the victim program does not leak secrets under straight-line speculation. More details and examples on how to verify a program’s compliance can be found in Section 5. We observe significant performance improvement when using Burst on our small code snippets. Interestingly enough, we also observe that (insecurely) activating burst mode on the whole enclave code leads to disastrous performance as it completely pollutes the TLB. More results and details are presented in Section 7.1. This highlights how Burst mode is complementary to SpecSafe and the importance of software-exposed controls for execution modes.
5. Security Analysis
Figure 7. Speculative control flow of memcpy assembly under our two mechanisms for secure shared memory
Citadel builds upon MI6, which already enforces Strong Microarchitectural Isolation (SMI) through partitioning and flushing (See Section 4.1 and (bourgeat2019mi6, ) for details). As a result, our security analysis focuses on the novelties brought forward by Citadel, i.e., shared memory between an enclave and the OS. This change effectively opens side channels in the shared memory hierarchy through the shared LLC partitions. Note that any other side channels (e.g., branch predictors) are still closed per inherited MI6’s isolation mechanisms: partitioning and flushing. Furthermore, we only worry about the confidentiality of the enclave with regard to the OS here (see Section 3). As explained in Section 3, the high-level security guarantee we are trying to enforce is Speculative Non-Interference (guarnieri2020spectector, ) (SNI) between the attacker and the victim program. To formally reason about SNI and capture these new interactions, we need a different set of arguments than for SMI and to leverage the framework of hardware software contracts for security introduced in Section 2.4. Following this methodology, we first define the security contracts we will consider.
5.1. Defining Security Contracts
A security contract essentially defines what architectural state of the victim is considered exposed or public. The contract is defined using two building blocks: the observer mode, or what architectural state is exposed, and the execution mode or which paths of a given program are considered (speculatively) reachable and exposed. We define an observer mode that views the sequence of addresses of every (speculative) victim access to shared memory, and the data of non-speculative accesses, but no other victim architectural state (i.e., the enclave PC is not exposed). We also consider that the sequence does not contain any timing information (i.e., no timestamps, only data and addresses). We call this observation mode “shared memory” abbreviated to shm.
We now describe the two different execution modes we consider for our contracts. The first execution mode we consider is the sequential mode (no speculation at all, i.e., one-instruction-at-a-time and in-order) or seq, which means that only non-speculative execution paths are considered reachable and addresses and data of shared memory accesses are exposed. We call this first contract seq_shm. Our second execution mode is the straight-line speculation mode or str, which means that we extend the sequential mode to also consider instructions up to a certain depth after a change in control flow (e.g., a branch or a jump) to expose their corresponding observations. We call this second contract str_shm.
5.2. Hardware Side
We now want to show that each of our hardware mechanisms satisfy one of these contracts, i.e., that they do not leak more information to an attacker that what is stated in the contract. In our case, we start from a design that satisfies Strong Microarchitectural Isolation. We also make sure to eliminate on-core microarchitectural state that could entangle state related to enclave private memory, shared memory and the attacker (see Section 6 where we bypass the L1-D cache). This means that the attacker is only able to observe the microarchitecture we explicitly share, i.e., the shared LLC partitions. As a result, the only architectural states the attacker is able to reconstruct through microarchitectural observations are the addresses of speculative shared memory accesses. Note that per our threat model described in Section 3 and following previous work, we artificially make the assumption that the attacker is not able to collect timestamps for the accesses it observes. For our first mechanism, SpecSafe (see Section 4.4), the program is executed following the out-of-order semantics of our processor, but we make sure to not touch any shared microarchitecture until a shared memory access has reached the head of the ROB and so is non-speculative. As a result, Citadel with SpecSafe verifies seq_shm. Our second mechanism is the burst mode (see Section 4.5). Here, the hardware was precisely designed to implement (see Section 4.5) straight-line speculation, and as a result obeys str_shm.
Listing 1: Pathological Insecure Code for Burst Mode
jal far_away_function
lw a0, 0(secret_address)
Listing 2: Assembly Snippet for the SHA256 library
csrwi 0x802,BURST
.loop1:
[...] # Straight-line code with
shared memory accesses
bne a3,a6,.loop1
csrwi 0x802,NO_BURST
5.3. Software Side
We now need to analyze if our programs can enforce SNI with regard to these contracts.
In our case, SNI reduces to the fact that the program does not expose more secret data, when considered under a given security contract than when run sequentially, i.e., under seq_shm. This is straightforward for our first contract, as the contract is exactly seq_shm. Putting everything together, this means that any enclave program secure when run sequentially (i.e., that does not explicitly leak its secrets) is also secure when run in Citadel with SpecSafe.
This is not true for all program for our second contract str_shm. Indeed, we can craft a pathological code snippet (see Listing 1) that would systematically leak a secret under str_shm while secure when run sequentially. Hence, burst mode can be harmful, and it is the responsibility of the programmer to make sure that they use burst mode in a safe context, by analyzing the code under the simple straight-line speculation model. Fortunately, we only use burst mode on small code snippets and are able to carry out the required program analysis by hand. This is possible as burst mode is enabled and disabled by writing to a special CSR with a csrwi instruction. This instruction cannot be performed speculatively and acts as a speculation barrier, hence it can be used to constrain the speculative flow control within clear bounds. Similar to previous work (guarnieri2021hardware, ; guarnieri2020spectector, ) we also make the assumption that when entering burst mode, the initial state of the microarchitecture touched by the code is independent of any secrets. This is reasonable as csrwi already flushes the processor pipeline, predictors are deactivated and the TLB and caches are easy to selectively warm up. The rest is left to the responsibility of the code before entering Burst mode.
SHA256 Example We annotate a LOAD macro in a SHA256 library and show the corresponding assembly snippet in Listing 2 (full assembly in supplemental material). The performance impact of this simple annotation can be seen in Section 7.1. Here, the assembly shows us that the code is enclosed between two csrwi, which bound speculation. Furthermore, one branch is present, but is immediately fenced by a csrwi. It is fenced as csrwi is a system instruction that cannot enter the reorder buffer until the reorder buffer is empty. As a result, for that specific code snippet, the speculative control flow, and as a result, the exposed architectural states are the same for str_shm as for seq_shm, which is sufficient to say that the code snippet realizes SNI with regard to str_shm.
memcpy Example We minimally annotate an implementation of memcpy to take advantage of burst mode when copying data from shared memory to enclave private memory (full C code in in supplemental material). More complex optimizations could use loop unrolling and annotate the second while loop. Performance results can be seen in Section 7.1. We show the details of the assembly for the code snippet we annotate in Figure 7. Here, the program analysis to verify SNI under str_shm is more complex as some shared memory accesses are performed speculatively. Nevertheless, by our definition of SNI, we only care about programs and inputs that lead to a secure execution (i.e., that does not leak secrets) under the seq_shm contract. That means that we can assume that registers a4 and a5 never hold any secret values under seq_shm as they are exposed to the attacker per our contract, through the shared memory accesses. Furthermore, at most one more run of .loop2 is executed speculatively. Hence, the only new values an attacker can learn are \$a4+8 and \$a5+8 with \$a4 and \$a5 the last values taken by the two registers under seq_shm and these values are not secret. Finally, which of these two values are exposed is public as we assume the initial microarchitectural state touched by the program to be public and no secrets are touched.
Program Analysis Tools Finally, for more complex pieces of code, static analysis could be done manually or using many existing tools: e.g.. oo7 (wang2019oo7, ), Kasper (johannesmeyer2022kasper, ), Kleespectre (wang2020kleespectre, ), Spectector (guarnieri2020spectector, ). Note that the usual critique and limitations of these tools are that 1) most code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model. Because Citadel also leverages hardware isolation, both of those limitations are not concerns. The code snippets we analyze are small and the observation and execution modes are simple, well-defined, and enforced by the hardware. Hence, these works are well-suited to our needs. We leave the automated analysis of programs by adapting these tools in the context of Citadel as future work (i.e., to support RISCV assembly).
6. A few technical details
We now elaborate on a few technical issues we faced while prototyping our machine. We also give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
Speculative Transmitters in the L1 D Cache When enabling shared memory, the core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses. While shared memory cannot be accessed speculatively, trusted memory can. It means that shared memory can be speculatively evicted by a speculative trusted memory access. Furthermore, this eviction will be observable to an attacker. The cache coherency protocol is such that writing to a memory location that is cached by another core will take more time than writing to a memory location that is not cached by another core. As a result, a shared L1 D can be leveraged by an attacker to create a speculative transmitter and mount a Spectre-style attack. Partitioning the L1-D cache would have a significant performance impact. Simply placing an arbiter in front of the L1-D arrays (which is very latency-sensitive) would have a much harsher overall impact on performance than for the LLC. At the same time, accesses to shared memory are typically not expected to be temporally reused (see Section 7.1), limiting the utility of the L1-D for shared memory anyway. For all these reasons, we decide to simply bypass the L1 D cache for enclave accesses to shared memory. Because the L1 is the only core-private microarchitectural state entangled with the adversary state through a coherence protocol, we do not need to worry about other core-private microarchitectural structures.
Miscellaneous Security Specifics We statically partition MSHRs between the two cores, implement a fair round-robin arbiter at the entry of the LLC (not implemented in (bourgeat2019mi6, )), and arbiter the TLB requests (ignored and omitted in (bourgeat2019mi6, )). Because these mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact in Citadel for insecure applications.
Reconfigurable LLC Partitioning An obvious bottleneck in our design is the static cache partitioning inherited from MI6. Dividing our 1MB LLC in 64 equal-size 16KB partitions (see Section 4.1) is a terrible idea as these are smaller than the 64KB L1 D+I caches. Because our caches are inclusive, an enclave or a program that resides in one memory region cannot even fully use the L1 caches. MI6 circumvented that issue by assuming and simulating (but not fitting on FPGA) a 16-core design with a larger 16MB LLC. We build a mechanism to reconfigure the LLC partitions, improving enclave and other programs’ performance, regardless of the number of cores or the size of the LLC. This costly operation is only required before launching a new application or enclave with a different memory requirement.
At the hardware level, we use a fine-grain set-partitioning scheme similar to (townley2022composable, ; saileshwar2021bespoke, ) to support our 64 partitions. Because for a given memory access, the memory region IDs are directly accessible through physical memory address, our scheme requires minimal changes in the LLC, no changes to the rest of the memory hierachy and supports shared memory between security domains. We also implement a fine-grain flushing mechanism by adding a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but which always sends back zeroes. This means that flushing a small LLC partition is much faster than flushing a big one. At the software level, the OS can request a reconfiguration through a SM call, which will ensure the LLC is not put in an incoherent state. The main challenge is to ensure that no core will emit requests to the modified regions during the operation. For the operation to proceed, the second core will have to perform a SM call to enter a loop located in a region unaffected by the reconfiguration. This means that an LLC reconfiguration is not possible if an enclave is running, some SM regions cannot be affected by the reconfiguration (checked by the SM). The SM will also check that the mapping is correct and fits within the LLC.
Building the Software Infrastructure The overall software architecture and privilege hierarchy can be seen in Figure 1. We implement the SM at the highest privilege level of RISC-V using a mix of assembler and C. We also implement what we call a Mini-SM, which is a downsized copy of the global SM private to each enclave. Second, we develop a Linux Kernel module that allows user programs to request SM calls and set up enclaves, keeping Linux responsible for the resource management. Finally, we implement a Secure Bootloader and an end-to-end attestation mechanism that makes it possible to uniquely identify an enclave, establish a shared secret with it and exchange encrypted private data.
@@ 7. Evaluation
Experimental Setup
Frontend 2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
Reorder Buffer 64 entries, 2-way insert/commit
Reservation Stations 2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
L1 (I/D) 32KB, 8 ways, 8 outstanding requests
Load-Store Unit 24 LdQ entries, 14 StQ entries, 4 SB entries
LLC (L2) 1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
Memory 120 cycles latency (24 outstanding requests)
@@
Citadel’s processor is a modified out-of-order 2-core based on Riscy-OOO (zhang2018composable, ) and MI6 (bourgeat2019mi6, ). The parameters used for the cores and the memory subsystem are given in Figure 8, and kept small to keep the synthesis and simulation time small. Citadel runs on an AWS EC2 F1 instance, equipped with a Virtex UltraScale+ FPGA. On this FPGA board, Citadel can be clocked at $30$ MHz, which is similar to the baseline design. We use the RISC-V tool chain with GCC 12.2.0. We compile all software with the -O3 flag. We run Linux Kernel 6.2.0. All performance numbers are collected by running the benchmarks on the FPGA using performance counters and an average of ten runs.
Implementation Data and TCB Breakdown
Component LOC Size
Processor (bourgeat2019mi6, ) (+ modification) $\sim 60$ KLOC (+ 250 ) N/A
Bootloader 5236 55KB
Security Monitor 9287 97KB
Mini-SM 2301 5.6KB
Linux $\sim 20$ MLOC 157MB
SM Kernel Mod. 507 505KB
TCB Software 14523 152KB
@@ 7.1. Evaluating Secure Shared Memory
Microbenchmark on memcpy We evaluate two different setups, one with the function as is (only using SpecSafe) and the other one when minimally annotated for Burst mode (full C code in supplemental material and assembly snippet in Figure 7). We normalize the results to using memcpy to and from private memory (not using our mechanisms) in Table 7.1. We observe that out-of-the-box, SpecSafe shows dramatic overhead, especially when the data is not cached in the LLC with a 2231% overhead. On the other end, we show that Burst mode is quite efficient and reduces the overhead to 245% and 62% depending on whether the data is cached in the LLC. We also try different configurations where we start with a cold LLC (all loads will go to main memory), start with a warm LLC (the OS has accessed it previously or the data just arrived from the network for instance) or an insecure setup, where we have enabled the L1 for shared memory accesses. Our results confirm that access latency has a huge impact on the performance of SpecSafe, which is mitigated by Burst Mode. Not shown in the table, we observe that when moving data from private memory to shared memory, the overhead is 48% in all cases (both mechanisms and with cold/warm LLC). Indeed, store operations to shared memory are already pipelined with SpecSafe and Burst mode does not improve the performance. We also measure the overhead of straight-line speculation to be around 33% by enabling Burst mode while using memcpy to and from private memory. Finally, looking at performance counters, we observe that the  20% overhead left can be attributed to address translations, which are already cached when performing private memory copy.
L1-D Bypass These results also highlight the reasonable performance impact of bypassing the L1-D for shared memory accesses, especially when using Burst mode with a warmed up LLC. Note that we can also make a qualitative argument for bypassing the L1-D as shared memory accesses usually do not involve re-access (computing on shared memory could be insecure as the access pattern is leaked) and getting the output data to the LLC as fast as possible makes a lot of sense as we want it to be accessed by another security domain.
SpecSafe Burst Mode
No LLC warm-up 23.317 3.458
LLC warm-up 3.665 1.621
LLC warm-up + L1 2.466 1.556
Table 2. memcpy micro-benchmark. Copy of 1500 bytes from shared memory to private memory. Normalized over memcpy from an to private memory with speculation enabled. We experiment with warming up the LLC and enabling the L1 for shared memory accesses (insecure).
{NiceTabular}
@@
Burst Only 1.654 4.959